# Can Uncertainty Quantification Enable Better Learning-based Index Tuning?

### v1

@@ 6.2 Results on Index Tuning
Figure 11 illustrates the index tuning results of nine methods under various budgets across six benchmarks. Table VI presents detailed statistics of cost improvements, highlighting the best performances in green and the second best in blue. From these results, we can draw the following conclusions:
First, integrating the UQ module into the two base models significantly improved index tuning results. The LIB model has seen a significant reduction in the number of worst-case scenarios, while the AMA model has experienced a substantial increase in the number of optimal configurations. Additionally, when integrating AE and Hybrid into the AMA model, the number of worst-case scenarios decreased from six to zero.
Second, the performance of the integrated Hybrid model is superior to that of the Ensemble-based model. As shown in Table III, Hybrid’s ability to quantify uncertainty surpasses that of Ensemble, allowing it to more accurately identify the parts of the learning-based BE model that it cannot accurately estimate, thereby improving BE’s accuracy. The DTA algorithm requires a specified maximum runtime, and usually, it cannot enumerate all index configurations by the end. Therefore, the greater the delay in BE-UQ inference, the fewer configurations the index advisor can explore. As seen in Table IV, Hybrid has a higher delay. Thus, by providing more accurate uncertainty quantifications, Hybrid finds better results with fewer configurations.
Third, AE’s results are slightly better than Hybrid’s. The accuracy of the LIB base model is not high, and in this situation, Hybrid’s UQ accuracy is higher than AE’s, enabling it to identify as many uncertain data points as possible, resulting in better performance than AE. Conversely, when using AMA as the base model, AE achieves higher identification accuracy and lower inference delay, allowing for more accurate enumeration of more index configurations. Therefore, it achieves the highest number of optimal configurations and surpasses the what-if tools. For index tuning tasks, Hybrid is preferable when there is sufficient time available for tuning. If time is limited and the BE model demonstrates high accuracy, selecting AE can result in a better index configuration.
Fourth, in scenarios abundant with Uncertainty Samples, the what-if approach remains a stable and precise method, achieving the best or nearly the best results across all six benchmarks. This consistency forms the foundation of our study, which aims to model the uncertainty in learning-based models, combining their accuracy and efficiency with the reliability of what-if tools.
7 Related Work
@@
We presented the related work on UQ in subsection 2.3 and subsection 5.2. The related work on learning-based BE models is detailed in subsection 2.2. In this section, we introduce additional related work on index tuning.
@@
Index Configuration Enumeration. Early greedy enumeration algorithms typically follow two strategies [38]. One is a top-down approach [21], such as the Drop algorithm [13], which iteratively removes candidate indexes with minimal impact on workload execution time until the budget constraint is met. The other is a bottom-up approach [14], exemplified by the Autoadmin algorithm proposed by Chaudhuri and Narasayya [12], which starts with an empty set and adds indexes until the maximum budget is reached. Recently, reinforcement learning-based methods, such as those by Lan et al. [11] and Sadri et al. [41], have shown greater adaptability and efficiency. SmartIX [42] introduces add/delete operations based on environmental state.
Workload Compression. To enhance the efficiency of index tuning, many approaches perform workload compression prior to tuning. Sadri et al. [41] identify templates based on the non-parameterized portions of query statements, retaining one query for each template. However, the query plans of a template can vary greatly depending on parameter differences, resulting in substantially different benefits calculated for the same index configuration and causing significant workload information loss. Furthermore, RIBE [39] clusters queries based on operators within the query plan, ensuring that only redundant information from the workload is eliminated. Besides these clustering-based methods, GSUM [2] and ISUM [43] define various query features and design greedy algorithms that flexibly determine the number of queries in the workload.
8 Conclusion
@@
In this paper, we address a critical yet unexplored question: Can uncertainty enable better learning-based index tuning? To answer this question, we focus on the core component of index tuning, prioritizing improvements to the learning-based BE model. Our investigation is divided into two sub-questions: (1) Can UQ methods accurately measure the reliability of model results? (2) How can a BE model that provides uncertainty estimates be integrated into an index advisor, and what are the implications of this integration?
@@
To address these two questions, we first analyze the origins and characteristics of model uncertainty in BE tasks. Unlike existing general UQ methods, we design a new hybrid quantification framework tailored to the task’s characteristics and propose an alternative version to enhance efficiency. Experimental validation shows that our method outperforms existing quantification methods in five out of six scenarios. The alternative version significantly improves UQ efficiency, albeit with a slight reduction in quantification accuracy. Next, we verified that the Beauty framework leverages the stability of what-if tools to supplement the learning-based BE model. In end-to-end experiments, this framework significantly improves the quality of index tuning results.
@@
Our findings confirm that integrating uncertainty can enhance learning-based index tuning. Our framework focuses on the BE problem, marking an initial exploration in this area. Future research will investigate the role of uncertainty in other index advisor modules.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (Grant No. 62072138).

### v2

@@ VI-B Results on Index Tuning
Figure 9 illustrates the index tuning results of nine methods under various budgets across six benchmarks. Table VII presents the ranking information of each benefit estimator across the six benchmarks, highlighting the best performances in green. Furthermore, Table VIII demonstrates the mean and variance of cost improvement achieved by each benefit estimator across every benchmark. From these results, we can draw the following conclusions:
Firstly, all three UQ methods enhance index tuning results compared to the base model: the frequency of worst-case scenarios decreases markedly, while the occurrence of optimal configurations increases. As evident from Table VIII, applying any UQ method to the base model yields improved cost performance, particularly with Hybrid and AE methodologies. This addresses the question: uncertainty quantification indeed enhances learning-based benefit estimation, resulting in superior index tuning outcomes.
Secondly, models incorporating the Hybrid approach demonstrate superior performance. Notably, AMA-Hybrid substantially outperforms other models in both cost improvement rankings and mean values. As illustrated in Table III, Hybrid’s capability to quantify uncertainty exceeds that of Ensemble and AE methods, enabling more precise identification of samples within the learning-based BE framework that cannot be accurately estimated, thereby enhancing BE’s precision. The DTA algorithm necessitates a predefined maximum runtime, and typically cannot enumerate all index configurations before termination. Consequently, greater delays in BE-UQ inference restrict the number of configurations the index advisor can explore. As evidenced in Table V, Hybrid exhibits increased latency. Nevertheless, by providing more accurate uncertainty quantifications, Hybrid achieves superior results with fewer configuration evaluations.
Thirdly, AE’s results surpass those of Ensemble. According to Table VIII, in $75\%$ of cases, the mean cost improvement when integrating AE exceeds that of Ensemble. This is further reflected in Table VII, wherein AE achieves significantly better rankings than Ensemble. These findings validate the effectiveness of our proposed UQ method based on the AutoEncoder architecture.
Fourthly, in scenarios replete with unreliable results, the what-if tools remain a stable and accurate methodology, significantly outperforming both base BE models. The effectiveness of these base BE models exhibits heightened sensitivity to benchmarks and storage constraints, thereby demonstrating insufficient robustness. This consistency forms the foundation of our investigation, which aims to model uncertainty in learning-based models, combining their computational efficiency and predictive accuracy with the reliability inherent in what-if tools.
VII Related Work
@@
We presented the related work on UQ in subsection II-C and subsection V-B. The related work on learning-based BE models is detailed in subsection II-B. This section focuses on the works concerning configuration enumeration algorithms.
@@
Early greedy enumeration algorithms typically follow two strategies [32]. One is a top-down approach [20], such as the Drop algorithm [12], which iteratively removes candidate indexes with minimal impact on workload execution time until the budget constraint is met. The other is a bottom-up approach [13], exemplified by the Autoadmin algorithm proposed by Chaudhuri and Narasayya [11], which starts with an empty set and adds indexes until the maximum budget is reached. Recently, reinforcement learning-based methods, such as those by Lan et al. [10] and Sadri et al. [36], have shown greater adaptability and efficiency. SmartIX [37] introduces add/delete operations based on environmental state.
VIII Conclusion
@@
In this paper, we address a critical yet unexplored question: Can uncertainty quantification enable better learning-based index tuning? To answer this question, we focus on the core component of index tuning by prioritizing improvements to the learning-based BE model. Our investigation revolves around two sub-questions: (1) Can UQ methods accurately measure the reliability of model results? (2) How can a BE-UQ model be integrated into an index advisor?
@@
To address these questions, we first analyze the origins and characteristics of model uncertainty in BE tasks. We then design a novel hybrid quantification framework and propose an efficient variant. Experiments show that our method outperforms existing methods in most cases. Additionally, we demonstrate that the Beauty framework significantly improves the quality of index tuning results.
@@
While our framework focuses on the BE problem, it represents an initial exploration in this area. Future research will investigate the role of uncertainty in other modules of the index advisor.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62072138, No. 62472123).